OpenStudy (anonymous):

what is the LCD of 55 and 45?

OpenStudy (dan815):

start from the lowest prime number and see if it goes into both numbers

OpenStudy (anonymous):

5*11 and 5*9 ?

OpenStudy (dan815):

yea

OpenStudy (anonymous):

so that would be 99?

OpenStudy (whpalmer4):

Factor both numbers into their prime factors: 45 = 3*3*5 55 = 5*11 Now multiply the highest power found of each unique factor 3*3*5*11

OpenStudy (anonymous):

so its 55? or would i multiply 55 with 3?

OpenStudy (whpalmer4):

3*3*5*11 =

OpenStudy (anonymous):

495

OpenStudy (whpalmer4):

That is correct. The lowest common multiple of 45 and 55 is 495. Multiples of 45: {45, 90, 135, 180, 225, 270, 315, 360, 405, 450, 495} Multiples of 55: {55, 110, 165, 220, 275, 330, 385, 440, 495, 550, 605} 495 is the first number that appears in both lists, agreed?

OpenStudy (anonymous):

oh alright thank you:)

OpenStudy (whpalmer4):

You can do the same procedure with polynomials, too.